Beam Test Meeting Mon. Apr. 23, 2012

From GlueXWiki
Jump to: navigation, search
  1. BCal Bias Scan (Andrei & Sascha)



- nothing has changed, still waiting on 5 pass photon beam

- took data at 5% of 2.5 GeV in pulst integral mode


- chiller died last night and temp fell to 2 C

- condensation on lines, but none observed inside detector

- see presetation above for data analysis done

- see elog for more details about daily activities

Long term plan is to move BCAL to 20 degree position next week, and to move FCAL to the 23% position in two weeks